OpenStudy (anonymous):

Trig: Find x so the distance between (x,2) and (1,5) is √13

OpenStudy (anonymous):

Distance formula!!

OpenStudy (anonymous):

How do you set it up?

OpenStudy (anonymous):

(x-2)^2 + (2-5)^2 = Root 13 I believe that is it

OpenStudy (anonymous):

x-1 **

OpenStudy (anonymous):

(1-x)^2+(5-2)^2=13 1+x^2-2x+9=13 1+x^2-2x-4=0 x^2-2x-3=0 solve for x x^2+1x-3x-3=0 x(x+1)-3(x+1)=0 (x+1)(x-3)=0 x=-1 0r 3

OpenStudy (anonymous):

Thank you Dhashni I wasn't sure if i was supposed to put the square root on top of the 13

OpenStudy (anonymous):

:)